OpenStudy (anonymous):

Simplify: 5^1c^2c^3c^4

OpenStudy (anonymous):

@mia99

OpenStudy (mia99):

sorry i have no idea what that is...... :(

OpenStudy (anonymous):

ok

OpenStudy (mia99):

is that like a quadratic equation??

OpenStudy (anonymous):

haha umm

OpenStudy (mia99):

\[5c ^{2}\times2c ^{3}\times4\] ??? is that what it look like

OpenStudy (anonymous):

nope

OpenStudy (mia99):

\[5^{1c ^{2c ^{3c ^{4}}}}\] Is that it??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

well kinda

OpenStudy (mathstudent55):

Is this the problem? \(\large 5^1c^2c^3c^4\)

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

What is a number to the 1 power equal to?

OpenStudy (mathstudent55):

Here's the rule: \(a^1 = a\)

OpenStudy (mathstudent55):

Based on the rule, what is 5^1 = ?

OpenStudy (anonymous):

uhh 5?

OpenStudy (mathstudent55):

Exactly. You already have the number part. Now we need the part with the c's.

OpenStudy (mathstudent55):

Here is the rule: \(\large a^m \times a^n = a^{m + n} \)

OpenStudy (mathstudent55):

You have \(c^2 \times c^3 \times c^4\) Write down c and add the exponents.

OpenStudy (mathstudent55):